Restart Night #12 UT 2018-09-13

OSA: J. Williams (LBTO) S. Allanson (Tucson)

Observers/ISA: B. Rothberg, J. Hill, (Tucson)

Software: S. Hooper (on-call), P. Kubanek, Y. Zhang (Tucson)

ISp: E. Solheid

Mtn Manager: M. Wagner

Details

UT 01:20 - GCS updates, fills complete, starting checklist to open, delayed handover due to J. Riedl's work in MODS

UT 01:45 - LUCI-2 MOS error on initialization
Content: Error to initialize the instrument unit server "rmi://localhost:60002/TWO_MOSUnit"!
Attachment: java.rmi.ServerException: RemoteException occurred in server thread; nested exception is:
java.rmi.RemoteException: de.rub.astro.lucifer.instrument.sequence.PreConditionsViolatedException
looks like strain gauge - mask 20 has been in fpu since last night, first thing i did was to initaliza all, GUI came up as UNKNOWN

-checked FPU to Storage in Mask Exchange on GUI - pre-conditions show both strain gauges on which is what would prevent it from going to storage

based on all of this - strain gauges on grabber are on and preventing initalize from completing

J. Power also checking on this and going through it with D. Thompson.

LBCs came on OK, 2nd night in a row

UT 02:09 - strain gauges zeroed, initialized ok

UT 02:19 - collimating now - will start with AGW-4 work

UT 02:26 ACTIVE preset to BS9178 on-axis. Both sides converge nicely, but DX pupil seems to be centered at 8,8 rather than 7,7 - both IDL and GCS agree. GCSL converged only to 1.3 arcsec FWHM in 0.9 arcsec seeing. Greg's new flat on SX AdSec looks a bit better near the center. This is the recently patched GCS which

2018-09-13T02:26:17.003596+00:00 tcs1 LBT_GCSL: CamImage: createSubImageXY(70 131 249 310)
2018-09-13T02:26:17.010574+00:00 tcs1 LBT_GCSL: [D] WFS: spot grid 0 0: [3:15,3:16]

UT 02:32 home AGW2 and resend the on-axis preset. DX pupil came back to the same place (perhaps only displaced by 0.75 lenslet rather than a full row).

UT 02:38 ACTIVE preset off-axis KFP X=-0.2 Y=69.74 SX FWHM=1.5 DX FWHM=0.8 arcsec.

2018-09-13T02:40:13.654168+00:00 tcs1 LBT_GCSL: CamImage: createSubImageXY(70 145 249 324)
2018-09-13T02:40:13.669151+00:00 tcs1 LBT_GCSL: [D] WFS: spot grid 0 0: [3:15,3:15]

UT 02:46 ACTIVE preset off-axis KFP X=0.1 Y=149.3 - DX grabbed the wrong guide star a fainter binary. (Interesting for seeing lenslet positions.) SX pupil shift is OK here on the guider Y-axis.

2018-09-13T02:48:46.684389+00:00 tcs1 LBT_GCSL: CamImage: createSubImageXY(70 158 249 337)
2018-09-13T02:48:46.696374+00:00 tcs1 LBT_GCSL: [D] WFS: spot grid 0 0: [3:15,3:16]

UT 02:55 resending preset to correct the guide star

UT 03:15 ACTIVE preset off-axis with PA rotated by +30 (after ACQUIRE for repointing) KFP X=+73.8 Y=129.8

grep GCSL /lbt/log/current.log | grep "createSubImageXY(" -A 2

Note that IDL and GCS on the SX side no longer agree on the left this is wfsc 0051. It must be a configuration problem as the DX side agrees.

SX
  => Pupil Positioning:              X       Y                   X       Y
        On-Axis Pupil =           159.70  214.76              159.70  214.76
          Pupil Shift =            17.58   22.26               17.58   22.26
 Shifted Pupil Center =           177.28  237.02
      Outer Pupil Fit =           177.67  236.89
                                  --------------              --------------
         Pupil Center =           173.11  245.11              173.07  234.77

DX
  => Pupil Positioning:              X       Y                   X       Y
        On-Axis Pupil =           376.00  263.00              376.00  263.00
          Pupil Shift =            17.58  -22.26               17.58  -22.26
 Shifted Pupil Center =           393.58  240.74
      Outer Pupil Fit =           396.31  247.75
                                  --------------              --------------
         Pupil Center =           397.99  245.20              397.99  245.20

03:22 ACTIVE preset off-axis with PA rotated by -30 from original. PA=115 SX side runs away immediately in coma! KFP X=-75.5 Y=128.8

SX wfsc 0060
  => Pupil Positioning:              X       Y                   X       Y
        On-Axis Pupil =           159.70  214.76              159.70  214.76
          Pupil Shift =            17.58   22.26               17.58   22.26
 Shifted Pupil Center =           177.28  237.02
      Outer Pupil Fit =           177.90  235.27
                                  --------------              --------------
         Pupil Center =           173.11  245.11              173.07  234.77

DX wfsc 0067
  => Pupil Positioning:              X       Y                   X       Y
        On-Axis Pupil =           376.00  263.00              376.00  263.00
          Pupil Shift =            17.58  -22.26               17.58  -22.26
 Shifted Pupil Center =           393.58  240.74
      Outer Pupil Fit =           394.13  246.68
                                  --------------              --------------
         Pupil Center =           397.99  245.20              397.99  245.20

after the preset

SX wfsc 0061
  => Pupil Positioning:              X       Y                   X       Y
        On-Axis Pupil =           159.70  214.76              159.70  214.76
          Pupil Shift =           -17.96   21.96              -17.96   21.96
 Shifted Pupil Center =           141.74  236.72
      Outer Pupil Fit =           145.04  235.93
                                  --------------              --------------
         Pupil Center =           146.35  231.73              153.00  215.00

DX wfsc 0068
  => Pupil Positioning:              X       Y                   X       Y
        On-Axis Pupil =           376.00  263.00              376.00  263.00
          Pupil Shift =           -17.96  -21.96              -17.96  -21.96
 Shifted Pupil Center =           358.04  241.04
      Outer Pupil Fit =           358.11  247.40
                                  --------------              --------------
         Pupil Center =           357.85  245.69              357.85  245.69

03:28 Clear ACTIVE optics on left, but image is still ugly.

03:33 ACTIVE preset on-axis

UT 03:50 restart GCS for new patch

UT 03:50 - sending preset to NGC 7027 LS 0.5"

UT 04:04 - acquiring star in NGC 7027UT

UT 04:07 - starting spectroscopy - star looks centered

UT 04:28 Doug: GCS lenslet grid could be offset in X so that one direction of pupil shift slides the pupil off the edge of the grid.

UT 04:38 - current guidestar was to the right of center in detector space, changing to a new guidestar left of center - will resend the preset (6/6 HK done, 4/6 zJ done, - redoing zJ only will do half the integration time)

-GCS stopped and restarted on both sides

UT 04:39 - Barry resends preset to -X guide star X=-55 Y=59 SX runs away, and subimage shift seems to be wrong.

UT 04:42 - looks like GS on SX is now having issues - image in guidecam (and later on images) are significantly poorer than before. Stopping zernikes, clearing active optics on both mirrors SX.

UT 04:50 change from 153 to 113 for hotspot_sext_x. that changes the lenslet index from -39 to 3. Petr will need to track down how -39 is even possible.

UT 04:51 send zernikes to PSF, and it does not run away, but collimation is still slightly triangular.

UT 04:55 change hotspot_sext_y from 214.7 to 224.3 that doesn't change the GCS calculated pupil center, or maybe it didn't take effect until the next preset. X=-40 Y=+76.5 Was a shift of sext_x by -40 enough? Doug sent better values the following day.

SX wfsc 220
  => Pupil Positioning:              X       Y                   X       Y
        On-Axis Pupil =           159.70  214.76              159.70  214.76
          Pupil Shift =            -7.41   14.14               -7.42   14.14
 Shifted Pupil Center =           152.29  228.90
      Outer Pupil Fit =           153.36  228.67
                                  --------------              --------------
         Pupil Center =           146.35  231.73              146.45  230.99

-image quality is so bad, can't see the central star on LUCI-1 but easily can see on LUCI-2

UT 04:58 -Going to Telluric (guidestar on the left side of det X)

LUCI-1 star not quite centered +0.2

LUCI-2 centered OK

UT 05:07 - telluric starting spectroscopy

UT 05:08 - error grating setup HK G200 1.93 on LUCI-1 - set manually -taking image manually - seemed to be ok

UT 05:24 - Imaging script to NGC 7027 - 5pt dither, GS on left, cycle through remaining filters, repeat of K, then skip to [FeII]

UT 05:25 accidental faint guide star shows lenslet field stops behind star.

UT 05:29 - resending preset, it selected a faint star (GS is bright, but LUCI script passed it R=15.8 mag) - new GS selected

UT 05:40 - skipping to [FeII] - filter error LUCI-2 - iniitalizing filter wheel one from IMGUI - came up as Y2 - manually moving to [FeII] - error moving to [FeII]

UT 05:42 - brought up LUCI-2 filter wheel unit from LMC - initialized - came up as Br Gamma - manually moving in steps of one filter - moved forward to clear, then back towards [FeII] - worked

UT 05:44 - (LUCI-2 will be missing first [FeII] exposure

UT 05:48 - image quality is about 20% worse on LUCI-1, but images are still round

UT 06:54 for LUCI_R.cfg installed Doug's new hotspot 376.0,263.0 ---> 377.5,265.5

UT 07:05 - completed NGC 7027 imaging script - have now taken data through all LUCI-filters on both side - last exposure luci1.20180913.0081, luci2.20180913.0079

Measuring wfs to LUCI focus

UT 07:06 - J. Hill stopping WFSing - going to test focus with LUCIs

/home/lbto/iraf/focus_luci.cl
# usage "cl < focus_luci.cl"

# Position 1

# move to the first focus position
!psfclient left instrument_offsets M2 abs 0 0 0.15 0 0 0

# offset the telescope positiom
offset 0 -10 type=ABS

!echo "Take LUCI image 1"

sleep 10
.........

UT 07:16 - going to a Persson star, LUCI-N 5sec Br Gamma measuring FWHM using LUCI RTDs

luci2.20180913.0081 0082 ignore

-restarting (turning off WFS)

UT 07:33 - 0083 start

83 14.11 16.87

84 11.93 10.07

85 5.78 6.81

86 5.01 4.49

87 7.45 7.95

88 12.95 8.49

89 17.34 9.72

#85 - was focus we were at already so all good

UT 07.40 repeat exercise LUCI-1 luci1.20180913.0082 first

82 12.33 7.51

83 6.64 6.77

84 5.08 4.06 - best

85 6.81 6.21

86 11.75 12.05

87 couldn't measure

-need to adjust focus

UT 07:53 Left focus from 31 to 26 mm, but that was the wrong way and 36 is correct (even though we used 37 for this test).

88 - test image to check focus 7.34 9.11 fwhm

-going other direction to focus=37

89 - 3.37 3.65 fwhm - BINGO!

LUCI image quality after focussing - NGC 628

UT 08:00 SX wfs says SATURATED On-Axis - see IT 7252

2018-09-13T08:14:53.903654+00:00 tcs1 LBT_GCSL: WFS: findSpots: mvSubImageXY (70, 154, 249, 333).
2018-09-13T08:14:53.903663+00:00 tcs1 LBT_GCSL: WFS: findSpots: mvSubImageXY (70, 154, 249, 333), shifted 0,0 pix.
2018-09-13T08:14:53.903671+00:00 tcs1 LBT_GCSL: WFS: findSpots: createSubImageXY (70, 154, 249, 333).
2018-09-13T08:14:54.062792+00:00 tcs1 LBT_GCSL: WFS: findSpots: GUI subImage filename = /Repository/GCSfiles/dailyimages/left_wfscsubimage000405.fits
2018-09-13T08:16:28.592000+00:00 tcs1 LBT_GCSL: WFS: findSpots: mvSubImageXY (23, -90, 202, 89).
2018-09-13T08:16:28.592006+00:00 tcs1 LBT_GCSL: WFS: findSpots: mvSubImageXY (23, 1, 202, 180), shifted 0,-91 pix.
2018-09-13T08:16:28.592017+00:00 tcs1 LBT_GCSL: WFS: findSpots: createSubImageXY (23, 1, 202, 180).
2018-09-13T08:16:28.749746+00:00 tcs1 LBT_GCSL: WFS: findSpots: GUI subImage filename = /Repository/GCSfiles/dailyimages/left_wfscsubimage000406.fits

UT 08:18 - sending preset to NGC 628 to test image quality after change - will test if pausing guiding for 300" will work

last exposure 0099 L1, 0098 L2

UT 08:53 changed SX WFScam_sext_x to 113.0,237.7

UT 09:04 Total collimation has NaNs. Cleared by initializing M2 LUTs in PSF, but they come back in a few minutes. Created IT 7404. Restarted PSF, GCS, PCS.

More LUCI/LUCI pointing data

UT 09:10 - going to take more pointing data

luci1.20180913.100, luci2.20180913.099,0100 - sky offs (2.51 x 2)
Sorted in Azimuth:

AZ.....ALT.......RA..J2000.......DEC............PtStarID......................L1............L2
======= ===================== ========
8.12 32.15 07:39:34.179 83:07:24.43 PT10m034......................0101......0101,0102
21.48 51.33 03:44:52.171 65:39:15.91 PT10m017......................0102.....0103,0104
48.18 66.54 03:11:04.380 45:58:09.12 PT10m015......................0103.....0105.0106
52.41 45.74 05:14:06.028 48:10:24.74 PT10m024......................0104.....0107,0108 - L2 set to NORMAL not integrated, fixed here
77.86 42.44 05:14:34.108 29:40:48.20 PT10m025......................0105.....0109
94.36 56.82 03:57:36.752 24:39:07.47 PT10m019......................0106.....0110
111.13 33.14 04:55:39.472 2:22:11.96 PT10m022......................0107.....0111
112.06 79.19 02:15:28.469 28:07:25.47 PT10m011......................0109...0113
126.94 60.96 03:03:54.241 13:06:35.48 PT10m014......................0110.....0114
142.14 37.61 03:28:55.041 -11:20:05.27 PT10m016......................0111.....0115
157.22 54.29 02:22:15.469 -0:48:41.75 PT10m012.......................0112....0116
167.47 35.89 02:13:14.566 -20:24:41.88 PT10m010......................0113.....0117
209.35 52.67 00:20:49.768 -0:52:15.96 PT10m002......................0114.....0118
212.96 77.26 01:00:18.152 21:47:27.61 PT10m005......................0115.....0119
234.69 59.62 23:49:52.999 12:42:56.46 PT10m102......................0116.....0120
271.04 61.81 23:19:22.350 28:54:26.91 PT10m100......................0117.....0121

UT 10:55 temp changed SX hotspot_y back to 218.6

UT 11:20 pointing check
314.29 62.19 23:28:13.610 48:46:01.30 PT10m101......................0118.....0122
347.05 77.12 01:13:44.415 45:11:40.71 PT10m008......................0119.....0123
351.44 49.98 00:19:22.047 71:35:43.13 PT10m001......................0120.....0124
42.42 46.64 07:05:30.801 55:01:05.09 PT10m032......................0121.....0125
7.50 35.67 07:39:34.179 83:07:24.43 PT10m034......................0122.....0126

21 stars
UT 09:21 - restarting all systems due to issue noted in IT 7407

UT 11:53 - putting LUCIs into safe config, closing up now, going to take LUCI cals for tonight's observations (why waste data that could be potentially useful to someone?)

UT 11:58 - powering off LBCs

UT 12:13 - another LUCI-1 grating error G200 1pt93 HKspec

UT 12:15 - staring LUCI-Bino G200 HKspec 0p5" slit 1.93 CW flats/arcs - took 1 manual 20sec dark after LUCI-N arcs

UT 12:35 - taking manual exposures for arcs for zJ (60sec, 5sec)

UT 12:46 - starting imaging flats

UT 13:30 - filter wheel 1 LUCI-2 had trouble getting to H2 - multiple failures keeps getting stuck at OH 1060 - stepping by 2s forward

UT 13:33 - finally got LUCI-2 FW1 to work

UT 13:45 - FINISHED


Topic revision: r4 - 14 Sep 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback